Calculate the inductance of the inductive sensor detecting a Carbon steel bar. (The Bar has an end on area of r x r )

μ0=4πx10-7H/m

n=63turns

t=3mm

r=2mm

d=2mm

R=36m

P=36m

µc IRON = 5000

µA Carbon steel = 100


permeability HA Air gap a R P ለ ለለNote: The core is not a semicircle you need to calculate the Reluctance due to the path length
Answer #1

Solution:

The inductance of the inductive sensor detecting a Carbon steel bar is calculated.

are: n = 63 turns raamm • Rivont esteedt Rair gap Solution Given values do= 4x1074/os ta 3mm d=2mm R=36m p=36m MC IRON & permRixon= Iiron MoMe Airon . Iiron PtartP Airon TOR Q TR2= TE2 .. Riron CP+ 28+P) MoMe (TT+2) 1 1 1 substituting these values, R& Rsteel a Car) MOMA ( (10) ..Rsteels 2136) 47TX 107 X 100 X TTX10-6 1823.78 x108 AT/Wb Rain l gap= MA ad ATTX 107 XA A= 975Perotal = Riront Rsteel thair gap Incluctance .L= na .H R Total (8.105 + 1823.78 + 1.125) x1 1.125) x108 A7 Wb =1833.01 X108